# Quantization of bounded symplectic domains associated with compact Lie groups

### v1

@@ 1 Introduction
Quantization of symplectic manifolds is a central and recurring theme in mathematical physics. Such manifolds naturally serve as phase spaces for classical mechanical systems, and their quantization provides a bridge from classical to quantum mechanics. The most familiar examples arise when the phase space is the cotangent bundle $T^{\ast}M$ of a configuration space $M$, whose topology may already be nontrivial. An even richer class of examples comes from constrained Hamiltonian dynamics, where the reduced (physical) phase spaces often exhibit nontrivial global geometry. The multitude of approaches to the quantization problem – commonly labeled as geometric, deformation, strict, Berezin–Toeplitz, etc. – makes any comprehensive review impossible here; we therefore refer the reader to the standard monographs [1, 2, 3, 4, 5].
The present work is motivated by a different class of examples, where the nontrivial topology of the phase space originates in momentum rather than configuration space. This shift of perspective introduces new features in the quantization problem and reveals novel connections between geometry and quantum theory. Historically, the hypothesis of curved momentum space was first put forward by Born [6, 7] and subsequently explored by many authors (see, e.g., [8, 9, 10, 11, 12] and references therein). One of the original motivations was the idea that a nontrivial momentum space geometry could act as an effective cut-off, mitigating or even eliminating ultraviolet divergences in quantum field theory. For instance, if momentum space is compact, then its finite volume provides a natural ultraviolet cut-off.
@@
Several mechanical systems on symplectic groupoids $T^{\ast}G\rightrightarrows\mathfrak{g}^{\ast}$ have been analyzed in recent works [25, 26]. These studies indicate that the compactness of momentum space can lead to notable physical consequences. For instance, the velocity of a free (nonrelativistic) particle is bounded from above, provided that the kinetic energy is a smooth function of the momenta. Furthermore, when a particle is subjected to a central potential, it cannot fall to the center, regardless of the strength of the attractive force. Conversely, for any repulsive central force, there exists a finite domain of bounded motion. These results challenge conventional intuition based on flat momentum spaces.
@@ 3.1 Classical observables
$\mathrm{Pol}^{n}(D)\subset\mathrm{Pol}^{n+1}(D)\,,\qquad\mathrm{Pol}^{n}(D)\cdot\mathrm{Pol}^{m}(D)\subset\mathrm{Pol}^{n+m}(D)\,$, (3.6)
@@
$\mathfrak{R}_{l_{1}}(G)\,\mathfrak{R}_{l_{2}}(G)_{l_{2}}\subset\mathfrak{R}_{l_{1}+l_{2}}(G)\,,\qquad\mathfrak{R}(G)=\bigcup_{l\geq 0}\mathfrak{R}_{l}(G)\,.$
Thus, the family $\mathfrak{R}(G)_{l}$ defines an exhaustive multiplicative filtration on $\mathfrak{R}(G)$, indexed by the abelian semigroup $(\mathbb{R}_{\geq 0},+)$.
@@ 3.2 Prequantization
$(f,h)=\int_{G}\overline{f(g)}h(g)dg\qquad\forall f,g\in\mathcal{H}\,$, (3.17)
@@
[…]
where the summation runs over all permutations of the symbols $x_{a_{1}},\ldots,x_{a_{n}}$. We will refer to $\mathcal{Q}_{\hbar}$ as the prequantization map. It follows immediately from the definition that
@@
for all $f,h\in\mathcal{F}(G)$. The first commutator shows that the operators $\mathcal{Q}_{\hbar}(x_{a})$ form a Lie algebra which is isomorphic to $\mathfrak{g}$ for $\hbar>0$. The second and third commutators then define a trivial extension of $\mathfrak{g}$ by the $\mathfrak{g}$-module $\mathcal{F}(G)$. Note that the action of $\mathfrak{g}$ on $\mathcal{F}(G)$ respects the filtration by the Casimir length, so that we can speak of an infinite-dimensional Lie algebra $\mathcal{L}=\mathfrak{g}\subset\hskip-10.00002pt+\mathcal{F}(G)$} filtered by the family of finite-dimensional subalgebras $\mathcal{L}_{l}=\mathfrak{g}\subset\hskip-10.00002pt+\mathcal{F}_{l}(G)$}. The prequantization map $\mathcal{Q}_{\hbar}$ defines a unitary representation of the universal enveloping algebra $\mathcal{U}(\mathcal{L})$ in the pre-Hilbert space $\mathcal{H}$. This allows us to identify the space of prequantum observables $\mathcal{Q}_{\hbar}(\mathfrak{A})$ with the quotient algebra $\mathcal{U}(\mathcal{L})/\mathcal{J}$, where the two-sided ideal $\mathcal{J}$ is generated by the relations
$\mathcal{Q}_{\hbar}(fh)-\mathcal{Q}_{\hbar}(f)\mathcal{Q}_{\hbar}(h)=0\,,\qquad\forall f,h\in\mathcal{F}(G)\,.$
@@
This shows that, when restricted to the subspace $\mathcal{H}^{l}$, the prequantization map is controlled by the $R$-norm, providing a concrete estimate of the operator size.
3.3 Quantazable semialgebraic domains
@@ 3.4 Quantization
The presence of the Laplace operator implies that $\dim\mathcal{H}^{D}\leq\infty$. The Weyl law (3.15) approximates the dimension of $\mathcal{H}^{D}$ for $R/\hbar\gg 1$. By abuse of language, we will refer to $\psi$ and $\mathcal{H}^{D}$ from (3.31) as the eigenstate and eigenspace of the domain $D$. If $D_{1}$ and $D_{2}$ are two quantizable basic domains that are compatible with each other, then $D_{1}\cup D_{1}$ is quantizable and we define
@@
The space $\mathcal{H}^{D}$ is identified with the state space of a quantum-mechanical system whose classical phase space is the $G\times D$. Since the factor $G$ represents the space of momenta, we can speak of a momentum representation for the space of states. Let $P_{D}$ denote the Hermitian projector onto the subspace $\mathcal{H}^{D}\subset\mathcal{H}$. By definition,

### v2

@@ 1 Introduction
Quantization of symplectic manifolds is a recurring theme in mathematical physics. Such manifolds naturally serve as phase spaces for classical mechanical systems, and their quantization provides a bridge from classical to quantum mechanics. The most familiar examples arise when the phase space is the cotangent bundle $T^{\ast}M$ of a configuration space $M$, whose topology may already be nontrivial. An even richer class of examples comes from constrained Hamiltonian dynamics, where the reduced (physical) phase spaces often exhibit nontrivial global geometry. The multitude of approaches to the quantization problem – commonly labeled as geometric, deformation, strict, Berezin–Toeplitz, etc. – makes any comprehensive review impossible here; we therefore refer the reader to the standard monographs [1, 2, 3, 4, 5].
The present work is motivated by a different class of examples, where the nontrivial topology of the phase space originates in momentum rather than configuration space. This shift of perspective introduces new features in the quantization problem and reveals novel connections between geometry and quantum theory. Historically, the hypothesis of curved momentum space was first put forward by Born [6, 7] and subsequently explored by many authors (see, e.g., [8, 9, 10, 11, 12, 13] and references therein). One of the original motivations was the idea that a nontrivial momentum space geometry could act as an effective cut-off, mitigating or even eliminating ultraviolet divergences in quantum field theory. For instance, if momentum space is compact, then its finite volume provides a natural ultraviolet cut-off.
@@
Several mechanical systems on symplectic groupoids $T^{\ast}G\rightrightarrows\mathfrak{g}^{\ast}$ have been analyzed in recent years [26, 27, 28, 29]. These studies indicate that the compactness of momentum space can lead to notable physical consequences. For instance, the velocity of a free (nonrelativistic) particle is bounded from above, provided that the kinetic energy is a smooth function of the momenta. Furthermore, when a particle is subjected to a central potential, it cannot fall to the center, regardless of the strength of the attractive force. Conversely, for any repulsive central force, there exists a finite domain of bounded motion. These results challenge conventional intuition based on flat momentum spaces.
@@ 3.1 Classical observables
$\mathrm{Pol}^{n}(D)\subset\mathrm{Pol}^{n+1}(D)\,,\qquad\mathrm{Pol}^{n}(D)\,\mathrm{Pol}^{m}(D)\subset\mathrm{Pol}^{n+m}(D)\,$, (3.6)
@@
$\mathfrak{R}_{l_{1}}(G)\,\mathfrak{R}_{l_{2}}(G)\subset\mathfrak{R}_{l_{1}+l_{2}}(G)\,,\qquad\mathfrak{R}(G)=\bigcup_{l\geq 0}\mathfrak{R}_{l}(G)\,.$
Thus, the family $\mathfrak{R}_{l}(G)$ defines an exhaustive multiplicative filtration on $\mathfrak{R}(G)$, indexed by the abelian semigroup $(\mathbb{R}_{\geq 0},+)$.
@@ 3.2 Prequantization
$(f,h)=\int_{G}\overline{f(g)}h(g)dg\qquad\forall f,h\in\mathcal{H}\,$, (3.17)
@@
$\begin{array}[]{c}\mathcal{Q}_{\hbar}(f)\psi=f\cdot\psi\,,\qquad\mathcal{Q}_{\hbar}(x_{a})\psi=-i\hbar\ell_{a}\psi\,,\qquad\forall f\in\mathcal{F}(G)\,,\quad\forall\psi\in\mathcal{H}\,,\\[14.22636pt] \displaystyle\mathcal{Q}_{\hbar}(fx_{a_{1}}\cdots x_{a_{n}})=\frac{1}{(n+1)!}\sum_{k=0}^{n}\sum_{\sigma\in S_{n}}\sigma\big(\mathcal{Q}_{\hbar}(x_{a_{1}})\cdots\mathcal{Q}_{\hbar}(x_{a_{k}})\mathcal{Q}_{\hbar}(f)\mathcal{Q}_{\hbar}(x_{a_{k+1}})\cdots\mathcal{Q}_{\hbar}(x_{a_{n}})\big)\,.\end{array}$ (3.18)
where the inner summation runs over all permutations of the symbols $x_{a_{1}},\ldots,x_{a_{n}}$. We will refer to $\mathcal{Q}_{\hbar}$ as the prequantization map. It follows immediately from the definition that
@@
for all $f,h\in\mathcal{F}(G)$. The first commutator shows that the operators $\mathcal{Q}_{\hbar}(x_{a})$ form a Lie algebra which is isomorphic to $\mathfrak{g}$ for $\hbar>0$. Then the second and third commutators generate a trivial extension of $\mathfrak{g}$ by the $\mathfrak{g}$-module $\mathcal{F}(G)$. Observe that the action of $\mathfrak{g}$ on $\mathcal{F}(G)$ respects the filtration by the Casimir length, so that we can speak of an infinite-dimensional Lie algebra $\mathcal{L}=\mathfrak{g}\subset\hskip-10.00002pt+\mathcal{F}(G)$} filtered by the family of finite-dimensional subalgebras $\mathcal{L}_{l}=\mathfrak{g}\subset\hskip-10.00002pt+\mathcal{F}_{l}(G)$}. The prequantization map $\mathcal{Q}_{\hbar}$ therefore defines a unitary representation of the universal enveloping algebra $\mathcal{U}(\mathcal{L})$ in the pre-Hilbert space $\mathcal{H}$. This allows us to identify the space of prequantum observables $\mathcal{Q}_{\hbar}(\mathfrak{A})$ with the quotient algebra $\mathcal{U}(\mathcal{L})/\mathcal{J}$, where the two-sided ideal $\mathcal{J}$ is generated by the relations
$fh-f\otimes h=0\,,\qquad\forall f,h\in\mathcal{F}(G)\,.$
Another possible interpretation of the commutation relations (3.19) is provided by the notion of Lie–Rinehart pair [37]. The Poisson algebra of classical observables $\mathfrak{A}$ contains the subspace $\mathfrak{A}^{0}\oplus\mathfrak{A}^{1}$ consisting of functions that are at most linear in $x_{a}$. The latter carries a natural structure of the Lie–Rinehart pair $(\mathfrak{M},\mathfrak{L})$, where the (infinite-dimensional) Lie algebra $\mathfrak{L}$ is given by $\mathfrak{A}^{1}$, and the $\mathfrak{L}$-module $\mathfrak{M}$ is identified with $\mathfrak{A}^{0}$. Both the Lie bracket on $\mathfrak{L}$ and the (left) action of $\mathfrak{L}$ on $\mathfrak{M}$ are defined by the Poisson bracket, while the left action of $\mathfrak{M}$ on $\mathfrak{L}$ is given by the dot product of functions. In geometrical terms, this Lie–Rinehart pair corresponds to the tangent Lie algebroid of $G$. With this interpretation, the quantization map $\mathcal{Q}_{\hbar}$ defines a homomorphism from the universal enveloping algebra $\mathcal{U}(\mathfrak{M},\mathfrak{L})$ of the Lie–Rinehart pair to the algebra of differential operators on $G$.
@@
This shows that when restricted to the subspace $\mathcal{H}^{l}$, the prequantization map is controlled by the $R$-norm, providing a concrete estimate of the operator size.
3.3 Quantizable semialgebraic domains
@@ 3.4 Quantization
The presence of the Laplace operator implies that $\dim\mathcal{H}^{D}<\infty$. The Weyl law (3.15) approximates the dimension of $\mathcal{H}^{D}$ for $R/\hbar\gg 1$. By abuse of language, we will refer to $\psi$ and $\mathcal{H}^{D}$ from (3.31) as the eigenstate and eigenspace of the domain $D$. If $D_{1}$ and $D_{2}$ are two quantizable basic domains that are compatible with each other, then $D_{1}\cup D_{1}$ is quantizable and we define
@@
The space $\mathcal{H}^{D}$ is identified with the state space of a quantum-mechanical system whose classical phase space is $G\times D$. Since the factor $G$ represents the space of momenta, we can speak of a momentum representation for the space of states. Let $P_{D}$ denote the Hermitian projector onto the subspace $\mathcal{H}^{D}\subset\mathcal{H}$. By definition,